The stockholders’ equity accounts of Pina Company have the following balances on December 31, 2017. Common stock, $10 par, 293,000 shares issued and outstanding $2,930,000 Paid-in capital in excess of par—common stock 1,190,000 Retained earnings 5,310,000 Shares of Pina Company stock are currently selling on the Midwest Stock Exchange at $33. Prepare the appropriate journal entries for each of the following cases. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.) (a) A stock dividend of 5% is (1) declared and (2) issued. (b) A stock dividend of 100% is (1) declared and (2) issued. (c) A 2-for-1 stock split is (1) declared and (2) issued. No. Account Titles and Explanation Debit Credit (a) (1) enter an account title for case A to record the declaration of stock dividends enter a debit amount enter a credit amount enter an account title for case A to record the declaration of stock dividends enter a debit amount enter a credit amount enter an account title for case A to record the declaration of stock dividends enter a debit amount enter a credit amount (a) (2) enter an account title for case A to record the issuance of stock dividends enter a debit amount enter a credit amount enter an account title for case A to record the issuance of stock dividends enter a debit amount enter a credit amount (b) (1) enter an account title for case B to record the declaration of stock dividends enter a debit amount enter a credit amount enter an account title for case B to record the declaration of stock dividends enter a debit amount enter a credit amount (b) (2) enter an account title for case B to record the issuance of stock dividends enter a debit amount enter a credit amount enter an account title for case B to record the issuance of stock dividends enter a debit amount enter a credit amount (c) (1) enter an account title for case C to record the declaration of the stock split enter a debit amount enter a credit amount enter an account title for case C to record the declaration of the stock split enter a debit amount enter a credit amount (c) (2) enter an account title for case C to record the issuance of the stock split enter a debit amount enter a credit amount enter an account title for case C to record the issuance of the stock split enter a debit amount
Answer #1

1)Number of shares issued as stock dividend =Number of shares outstanding*% of stock dividend

                   = 293000*5%

                   = 14650 shares

Date Account title Debit credit
a Retained earning (14650*33) 483450
common stock dividend Distributable (14650*10) 146500
Paid-in capital in excess of par—common stock 336950
b common stock dividend Distributable 146500
common stock 146500

2)

Number of shares issued as stock dividend =Number of shares outstanding*% of stock dividend

               = 293000*100%

                = 293000 shares.

Date Account title Debit credit
a Retained earning (293000*10) 2930000
common stock dividend Distributable 2930000
b common stock dividend Distributable 2930000
common stock 2930000

3)stock split does not require any adjustment entry .The only adjustment is that number of shares outstanding after stock split is increased and par value gets decreased.

Number of shares outstanding after stock split = 293000*2/1 = 586000 shares

par value per share = 10*1/2 = $ 5 per share

Date Account title Debit credit
a No entry
b No entry